The weight of a sophisticated running shoe is normally distributed with a mean of 12 ounces
and a standard deviation of 0.5 ounces.
a. What is the probability that a shoe weighs more than 13 ounces?
b. What must the standard deviation of weight be in order for the company to state that 99.9%
of its shoes are less than 13 ounces?

Answer by Boreal(15235)

z=(x-mean)/sd
=(13-12)/0.5=2 sd
the probability is z>2 or 0.0228
-
99.9% is a z-value of 3.09
so 3.09=(1)/sd
so sd has to be 1/3.09 or 0.324 oz.